a brick of weight 3000gf and size 25cm x 10cm x 6cm is placed in three different positions on the ground. Find the pressure exerted by the brick in each position on the ground.
1
Expert's answer
2021-09-13T12:16:34-0400

(a)

"P=\dfrac{F}{A}=\dfrac{mg}{a\cdot b},""P=\dfrac{3000\ gf\cdot\dfrac{9.8\cdot10^{-3}\ N}{1\ gf}}{0.25\ m\cdot 0.1\ m}=1176\ Pa."

(b)

"P=\dfrac{F}{A}=\dfrac{mg}{a\cdot c},""P=\dfrac{3000\ gf\cdot\dfrac{9.8\cdot10^{-3}\ N}{1\ gf}}{0.25\ m\cdot 0.06\ m}=1960\ Pa."

(c)

"P=\dfrac{F}{A}=\dfrac{mg}{b\cdot c},""P=\dfrac{3000\ gf\cdot\dfrac{9.8\cdot10^{-3}\ N}{1\ gf}}{0.1\ m\cdot 0.06\ m}=4900\ Pa."